OpenStudy (anonymous):

Simplify Radical 28/75

OpenStudy (mathstudent55):

This? \( \sqrt{\dfrac{28}{75} }\)

OpenStudy (anonymous):

Yeah

OpenStudy (mathstudent55):

Let's look at the prime factors of 28 and 75: \( \sqrt{\dfrac{2 \times 2 \times 7}{3 \times 5 \times 5} }\) This allows you to simplify this a bit before rationalizing.

OpenStudy (anonymous):

Yeah

OpenStudy (mathstudent55):

\( =\sqrt{ \dfrac{4}{25}} \sqrt{\dfrac{7}{3} }\) \(= \dfrac{2}{5} \sqrt{\dfrac{7}{3} }\) \(= \dfrac{2}{5} \dfrac{\sqrt{7}}{\sqrt{3}}\)

OpenStudy (mathstudent55):

There is still a sqrt(3) in the denominator. Now we need to rationalize the denominator.

OpenStudy (mathstudent55):

\(= \dfrac{2}{5} \dfrac{\sqrt{7}}{\sqrt{3}} \dfrac{\sqrt{3}}{\sqrt{3}}\) \(= \dfrac{2}{5} \dfrac{\sqrt{21}}{3} \) \(= \dfrac{2\sqrt{21} }{15} \)